giogang:

Two radio signals are broadcast. The radio waves can be illustrated as a hyperbola with a center at 0,0, a focus of 75,0, and a vertex of 56,0 Find the equation that describes the radio waves.

ZombieLover:

Hi! Welcome to Questioncove How may i help you??

surjithayer:

distance between center and focus=c=? distance between center and vertex=a=? \[c^2=a^2+b^2\] b=? eq. of hyperbola is \[\frac{ (x-h)^2 }{ a^2 }-\frac{ (y-k)^2 }{ b^2 }=1\] where (h,k) is center of hyperbola. h=0 k=0 can you complete it?